OpenStudy (jackiem272):

use separation of variables to solve the initial value problem dy/dt= y^2 and observe that the solution has a vertical asymptote at t=1. this phenomenon is called blow-up..because the solution becomes infinite in a finite time. but then it also asks to solve the initial problem using modified euler's method in excel

OpenStudy (solomonzelman):

where is your initial value?

OpenStudy (solomonzelman):

In general, if you have \(\color{#000000}{ \displaystyle \frac{dy}{dt} =y^n }\) You can bring the y-terms to the other side by division. \(\color{#000000}{ \displaystyle y^{-n}\frac{dy}{dt} =1 }\) then integrate both sides with respect to t. \(\color{#000000}{ \displaystyle \int y^{-n}\frac{dy}{dt} dt= \int 1 dt }\) Note that the dt cancels on the left side. \(\color{#000000}{ \displaystyle \int y^{-n}dy= \int dt }\) Then, for \(n\ne 1\), you get: \(\color{#000000}{ \displaystyle \frac{y^{-n+1}}{-n+1}=x+C }\) \(\color{#000000}{ \displaystyle y^{-n+1}=(-n+1)x+(-n+1)C }\) Number (-n+1) × arbitrary constant C is still arbitrary constant C, so, \(\color{#000000}{ \displaystyle y^{-n+1}=(-n+1)x+C }\) and you can write it explicitely if -n+1 is odd.

OpenStudy (solomonzelman):

then, if you are given that y(a)=b, then \(\color{#000000}{ \displaystyle a^{-n+1}=(-n+1)b+C }\) \(\color{#000000}{ \displaystyle a^{-n+1}-(-n+1)b=C }\) \(\color{#000000}{ \displaystyle a^{-n+1}+(n-1)b=C }\) So, \(\color{#000000}{ \displaystyle y^{-n+1}=(-n+1)x+a^{-n+1}+(n-1)b }\)

OpenStudy (solomonzelman):

but, this is all abstract of course. You would get a simpler result if you tried the same process with your equation.